OpenStudy (sagewilson):

Will medal Find the length of an arc of 40° in a circle with an 8 inch radius. 8π/9 inches <-- NOT correct. 16π/9 inches 64π/9 inches

OpenStudy (anonymous):

\[\frac{\pi 16\ 40}{360}=\frac{16 \pi }{9}=5.58505 \]

OpenStudy (sagewilson):

Oh okay, I see how you did it. Thank you for the help.

OpenStudy (anonymous):

Your welcome.